OpenStudy (anonymous):

on a map, the scale is 1 inch=2.5 miles. find the dimensions of a school district that is 4 inches by 9 inches on a map

Directrix (directrix):

1/(2.5) = 4/x x = 4 (2.5) x = 10 miles 1/(2.5) = 9/y y = 9(2.5) y = 22.5 miles Dimensions: 10 miles by 22.5 miles